# GraphLand: Evaluating Graph Machine Learning Models on Diverse Industrial Data

### v3

@@ 2 Limitations of popular graph machine learning benchmarks
By far the most popular datasets used in modern GML literature are the three academic citation networks cora, citeseer, and pubmed (Giles et al., 1998; McCallum et al., 2000; Sen et al., 2008; Namata et al., 2012; Yang et al., 2016). These datasets became so widespread perhaps because they were used by the foundational work on modern GNNs by Kipf & Welling (2017). However, these datasets only cover a single and rather narrow application of paper subject prediction in citation networks. Another popular set of datasets for GML was introduced by Shchur et al. (2018) and includes academic coauthorship networks coauthor-cs and coauthor-physics, and e-commerce co-purchasing networks amazon-computers and amazon-photo. However, all the aforementioned datasets together only cover three applications, while GML methods can be used in a much wider variety of settings. Later, larger-scale graph datasets were introduced in the Open Graph Benchmark (OGB) (Hu et al., 2020). However, out of the five node property prediction datasets, three are academic citation networks (ogbn-arxiv, ogbn-mag, ogbn-papers100M) and one more is an e-commerce co-purchasing network (ogbn-products). Thus, OGB does not significantly expand the range of real-world applications available for evaluating GML models. Further, all the aforementioned datasets only provide textual descriptions as node features. However, co-purchasing networks, which represent a very practically important application of GML, in realistic settings come with rich product metadata that can be represented as numerical and categorical features. There is currently a lack of datasets with such metadata available among popular GML benchmarks. Further, all the aforementioned datasets are homophilous, i.e., edges in them typically connect nodes of the same class. It has been shown by Huang et al. (2021) that even very simple models can provide strong results in homophilous networks. Thus, it is important for standard GML benchmarks to also include a wide selection of non-homophilous graphs, i.e., graphs in which edges do not have the tendency to connect nodes of the same class. For a long time, the only popular source of non-homophilous graph datasets was the benchmark from Pei et al. (2020). However, it was recently shown by Platonov et al. (2023b) that these datasets have numerous problems including duplicated nodes, small size (leading to noisy evaluation metric estimates), and insufficient class representation (such as the texas dataset having a class that consists of only a single node). While Platonov et al. (2023b) introduced several new non-homophilous graph datasets, they were meant to be used to reliably reevaluate the performance of different models in absence of homophily, rather than represent realistic GML applications. Thus, some of these datasets are synthetic (minesweeper), semi-synthetic (roman-empire), or have limited node features despite the original data source potentially providing more information about the nodes (tolokers, questions).
@@
Overall, it can be seen that the currently popular graph datasets for node property prediction do not allow evaluating GNNs and other GML models on a wide range of practically impactful industrial applications.
@@
Most of the datasets frequently used for node property prediction only have textual descriptions as node attributes. However, graphs representing real-world networks often have rich and diverse node attributes that go beyond just texts and encompass a variety of numerical and categorical features with different meanings and distributions. There has recently been a lot of interest in developing general-purpose GFMs that are expected to generalize to graphs from different domains (Wang et al., 2025; Mao et al., 2024). A key challenge for such GFMs is being able to adapt to graphs with different node feature sets which is required for a model truly generalizable to different domains. Yet, the GFMs that have been proposed in the current literature typically overlook this challenge and are often only evaluated on text-attributed graphs (Liu et al., 2024; Li et al., 2024a; He et al., 2025). Textual attributes can be easily projected to a common latent feature space by applying pretrained text encoders based on Large Language Models, thus allowing a single GFM to work with different text-attributed graphs. The prevalence of such text-attributed graphs in GML benchmarks has led to most of current GFM research overlooking the problem of generalization to different node feature sets, since it is not required to solve tasks from standard benchmarks. However, this problem is very important for real-world industrial applications of GML in which graphs often come attributed with a mixture of numerical and categorical node features. GFMs must therefore be able to work with such features to effectively solve practical tasks. The problem of devising a single foundation model that can work with arbitrary numerical and categorical features has received significant attention from the ML for tabular data community, where such features are standard, and first successful attempts to develop such a model have recently emerged (Hollmann et al., 2023, 2025; Ye et al., 2023, 2025; Mueller et al., 2025; Ma et al., 2024; QU et al., 2025). However, these ideas have not yet spread to the GML community (likely because current standard GML benchmarks do not require working with non-textual node features) despite the significant benefits of designing a successful GFM that can handle arbitrary node feature sets for practical applications.
@@ 3 GraphLand: a collection of diverse industrial graph datasets
  • If a graph is dynamically evolving in time, it should ideally have the necessary temporal information to construct a realistic time-based data split and a meaningful inductive setting;
@@
In this section, we briefly describe our datasets and the corresponding prediction tasks. Much more detailed descriptions are provided in Appendix A.1, while the discussion of structural properties and other characteristics of our datasets can be found in Appendix A.2.
@@
web-traffic, web-fraud and web-topics   These three datasets represent a part of the Internet (web graph). The nodes are websites, and a directed edge connects two nodes if at least one user followed a link from one website to the other in a selected period of time. We prepared three datasets with the same graph but different tasks: in web-traffic, the task is to predict how many users visited a website in a specific period of time (regression); in web-fraud, the task is to predict which websites are fraudulent (strongly imbalanced binary classification); and in web-topics, the task is to predict the topic that a website belongs to (multiclass classification). With almost $3$ million nodes, this is one of the largest publicly available attributed graphs that is not a citation network.
artnet-views and artnet-exp   These two datasets represent a social network of art creators. The nodes are users, and a directed edge connects two nodes if one user is subscribed to the other. We prepared two datasets with the same graph but different tasks: in artnet-views, the task is to predict how many views a user gets in a specific period of time (regression); and in artnet-exp, the task is to predict which users tend to create explicit art content (binary classification).
@@
hm-categories and hm-prices   These two datasets are based on the product co-purchasing network from the H&M company (Garc´ıa Ling et al., 2022). The nodes are products, and an edge connects two nodes if the products are often bought by the same customers. We prepared two datasets with the same graph but different tasks: in hm-categories, the task is to predict the product category (multiclass classification); and in hm-prices, the task is to predict the product price (regression).
@@
twitch-views   This dataset is based on the data from Rozemberczki & Sarkar (2021). It represents the live-streaming network Twitch. The nodes are users, and a directed edge connects two nodes if one user is subscribed to the other. The task is to predict how many views a user receives in a specific period of time (regression).
@@
Overall, our datasets cover a diverse range of industrial applications, and, as we show in Appendix A.2, they exhibit very diverse graph structural properties.
@@ 4 Data splits and experimental settings
In GML literature, there are two most popular settings for node property prediction regarding the relative sizes of train, validation, and test sets: one with a high label rate and one with a low label rate. In the high label rate setting, the train set encompasses 5050% of all graph nodes or more. This setting is common in heterophilous benchmarks, e.g., it is used by datasets from Pei et al. (2020) and Platonov et al. (2023b), and it is also used by most datasets from OGB (Hu et al., 2020). In the low label rate setting, much smaller train set sizes are used (typically, no more than 1010% of all graph nodes). This setting is commonly used with the classic cora, citeseer, pubmed citation networks, and it is also used by the ogbn-products dataset from OGB. Both of these settings can appear in real-world GML usage scenarios, depending on the resources available for data labeling. It is important to provide predetermined splits to ensure experiments in different works are run in the same setting and their results are comparable, but it is also important to accommodate different needs of different research projects. Thus, for datasets in our benchmark we provide fixed splits for both settings. We refer to these splits as the RL (random low) and RH (random high) splits. Specifically, the RL split randomly divides nodes into train/validation/test sets with 1010%/1010%/8080% proportions, while the RH split randomly divides nodes into train/validation/test sets with 5050%/2525%/2525% proportions.
@@
The RL and RH data splits are random, as is common in current GML benchmarks. However, in real-world applications, data splits are often temporal, i.e., the labeled objects are the ones that appeared in the network earlier, while the ones that appeared later are not labeled and belong to the test set. Despite their prevalence in applications, temporal splits are very rarely used in current GML node property prediction benchmarks. To the best of our knowledge, the only such datasets with temporal splits available are citation networks from OGB, which represent only a single application. At the same time, temporal data splits may significantly affect the prediction problem and the model performance, as they often result in distributional shifts between train, validation, and test data. While some types of distributional shifts have been previously explored in the GML literature, e.g., shifts in node features (Gui et al., 2022) and shifts in graph structure (Bazhenov et al., 2023), realistic temporal distributional shifts often combine shifts in several aspects of data simultaneously (e.g., shifts in the distributions of node features, labels, and graph structural characteristics), and the effect of such realistic shifts on GML model performance is currently under-explored. To close this gap, we provide a temporal split for most datasets in our benchmark. We refer to this split as the TH (temporal high) split; it divides nodes into train/validation/test sets with 5050%/2525%/2525% proportions, i.e., exactly the same proportions as in the RH split, which allows comparing model results between the RH and TH splits to see how the complexity of the task changes when temporal distributional shifts are introduced.
@@
Further, many real-world networks are not static, but evolve over time. Thus, in many applications, not only are there no labels available for nodes that appear in the network later, but the nodes themselves (with their attributes and incident edges) are not available at training time. This setting is known in GML as the inductive setting. In contrast to the transductive setting in which the whole graph is available at training time (including the nodes for which predictions should be made), in the inductive setting validation and test nodes are not available at training time. Despite temporally evolving graphs being common in practical applications, most standard node property prediction datasets only provide the transductive setting. While it is well-known that GNNs, in contrast to some other GML methods like shallow node embeddings, can work not only in the transductive but also in the inductive setting (Hamilton et al., 2017), it is not well-explored how the lack of complete graph information at training time in the inductive setting affects GNN performance. Moreover, when the inductive setting is used for GNN evaluation in the current literature, the validation and train nodes are typically chosen randomly, which is not realistic, since in real-world applications the inductive setting is almost always induced by the temporal evolution of the graph. To fill this gap, for all datasets in our benchmark for which temporal information is available, we additionally provide the inductive experimental setting. We refer to this setting as THI (temporal high / inductive); it has the exact same data split as the transductive TH setting, but provides three snapshots of the graph: one for training, one for validation, and one for testing. This allows investigating how model performance changes between the transductive and the inductive setting. To the best of our knowledge, our work is the first to compare GNN performance between random and temporal data splits in the transductive setting under the same split ratios, and between transductive and inductive settings under the same temporal data split. This comparison allows us to investigate how much these differences can affect GNN performance.

### v4

@@ 2 Limitations of popular graph machine learning benchmarks
By far the most popular datasets used in modern GML literature are the three academic citation networks cora, citeseer, and pubmed (Giles et al., 1998; McCallum et al., 2000; Sen et al., 2008; Namata et al., 2012; Yang et al., 2016). These datasets became so widespread perhaps because they were used by the foundational work on modern GNNs by Kipf & Welling (2017). However, these datasets only cover a single and rather narrow application of paper subject prediction in citation networks. Another popular set of datasets for GML was introduced by Shchur et al. (2018) and includes academic coauthorship networks coauthor-cs and coauthor-physics, and e-commerce co-purchasing networks amazon-computers and amazon-photo. However, all the aforementioned datasets together only cover three applications, while GML methods can be used in a much wider variety of settings. Later, larger-scale graph datasets were introduced in the Open Graph Benchmark (OGB) (Hu et al., 2020). However, out of the five node property prediction datasets, three are academic citation networks (ogbn-arxiv, ogbn-mag, ogbn-papers100M) and one more is an e-commerce co-purchasing network (ogbn-products). Thus, OGB does not significantly expand the range of real-world applications available for evaluating GML models. Further, all the aforementioned datasets only provide textual descriptions as node features. However, co-purchasing networks, which represent a very practically important application of GML, in realistic settings come with rich product metadata that can be represented as numerical and categorical features. Popular GML benchmarks currently lack datasets with such metadata. Further, all the aforementioned datasets are homophilous, i.e., edges in them typically connect nodes of the same class. It has been shown by Huang et al. (2021) that even very simple models can provide strong results in homophilous networks. Thus, it is important for standard GML benchmarks to also include a wide selection of non-homophilous graphs, i.e., graphs in which edges do not have the tendency to connect nodes of the same class. For a long time, the only popular source of non-homophilous graph datasets was the benchmark from Pei et al. (2020). However, it was recently shown by Platonov et al. (2023b) that these datasets have numerous problems including duplicated nodes, small size (leading to noisy evaluation metric estimates), and insufficient class representation (such as the texas dataset having a class that consists of only a single node). While Platonov et al. (2023b) introduced several new non-homophilous graph datasets, they were meant to be used to reliably reevaluate the performance of different models in the absence of homophily, rather than represent realistic GML applications. Thus, some of these datasets are synthetic (minesweeper), semi-synthetic (roman-empire), or have limited node features despite the original data source potentially providing more information about the nodes (tolokers, questions).
@@
Overall, the currently popular graph datasets for node property prediction do not allow evaluating GNNs and other GML models on a wide range of practically impactful industrial applications.
@@
Most of the datasets frequently used for node property prediction only have textual descriptions as node attributes. However, graphs representing real-world networks often have rich and diverse node attributes that go beyond texts and encompass a variety of numerical and categorical features with different meanings and distributions. There has recently been a lot of interest in developing general-purpose GFMs that are expected to generalize to graphs from different domains (Wang et al., 2025; Mao et al., 2024). A key challenge for such GFMs is being able to adapt to graphs with different node feature sets, which is required for a model truly generalizable to different domains. Yet, the GFMs that have been proposed in the current literature typically overlook this challenge and are often only evaluated on text-attributed graphs (Liu et al., 2024; Li et al., 2024a; He et al., 2025). Textual attributes can be easily projected to a common latent feature space by applying pretrained text encoders based on Large Language Models, thus allowing a single GFM to work with different text-attributed graphs. The prevalence of such text-attributed graphs in GML benchmarks has led to most of current GFM research overlooking the problem of generalization to different node feature sets, since it is not required to solve tasks from standard benchmarks. However, this problem is very important for real-world industrial applications of GML in which graphs often come attributed with a mixture of numerical and categorical node features. GFMs must therefore be able to work with such features to effectively solve practical tasks. The problem of devising a single foundation model that can work with arbitrary numerical and categorical features has received significant attention from the ML for tabular data community, where such features are standard, and first successful attempts to develop such a model have recently emerged (Hollmann et al., 2023, 2025; Ye et al., 2023, 2025; Mueller et al., 2025; Ma et al., 2024; Qu et al., 2025). However, these ideas have not yet spread to the GML community (likely because current standard GML benchmarks do not require working with non-textual node features) despite the significant benefits of designing a successful GFM that can handle arbitrary node feature sets for practical applications.
@@ 3 GraphLand: a collection of diverse industrial graph datasets
  • If a graph is dynamically evolving over time, it should ideally have the necessary temporal information to construct a realistic time-based data split and a meaningful inductive setting;
@@
In this section, we briefly describe our datasets and the corresponding prediction tasks. More detailed descriptions are provided in Appendix A.1, while the discussion of structural properties and other characteristics of our datasets can be found in Appendix A.2.
@@
web-fraud, web-topics, and web-traffic   These three datasets represent a part of the Internet (web graph). The nodes are websites, and a directed edge connects two nodes if at least one user followed a link from one website to the other in a selected period of time. We prepared three datasets with the same graph but different tasks: in web-fraud, the task is to predict which websites are fraudulent (strongly imbalanced binary classification); in web-topics, the task is to predict the topic that a website belongs to (multiclass classification); and in web-traffic, the task is to predict how many users visited a website in a specific period of time (regression). With almost $3$ million nodes, this is one of the largest publicly available attributed graphs that is not a citation network.
artnet-views and artnet-exp   These two datasets represent a social network of art creators. The nodes are users, and an edge connects two nodes if the users are friends. We prepared two datasets with the same graph but different tasks: in artnet-views, the task is to predict how many views a user receives in a specific period of time (regression); and in artnet-exp, the task is to predict which users tend to create explicit art content (binary classification).
@@
hm-categories and hm-prices   These two datasets are based on the product co-purchasing network from the H&M company (García Ling et al., 2022). The nodes are products, and an edge connects two nodes if the products are often bought by the same customers. We prepared two datasets with the same graph but different tasks: in hm-categories, the task is to predict the product category (multiclass classification); and in hm-prices, the task is to predict the product price (regression).
@@
twitch-views   This dataset is based on the data from Rozemberczki & Sarkar (2021). It represents the live-streaming network Twitch. The nodes are users, and an edge connects two nodes if both users follow each other. The task is to predict how many views a user receives in a specific period of time (regression).
@@
Overall, our datasets cover a diverse range of industrial applications. In Table 1, we provide some characteristics of the datasets: as can be seen, our datasets also exhibit very diverse graph structural properties. A more detailed discussion of these characteristics is provided in Appendix A.2.
Table 1: Characteristics of the proposed GraphLand datasets.
node classification node regression
hm-categories
pokec-regions
web-topics
tolokers-2
city-reviews
artnet-exp
web-fraud
hm-prices
avazu-ctr
city-roads-M
city-roads-L
twitch-views
artnet-views
web-traffic
median degree $45$ $13$ 22 $30$ 44 22 22 $45$ $71$ 44 33 $32$ 22 22
diameter $13$ $14$ $36$ $11$ 1919 $13$ $36$ $13$ $14$ $383$ $553$ 88 $13$ $36$
# classes 2121 $183$ $28$ 22 22 22 22 N/A N/A N/A N/A N/A N/A N/A
# node features $35$ $56$ $263$ $16$ $37$ $75$ $266$ $41$ $260$ $26$ $26$ 44 $50$ $267$
@@ 4 Data splits and experimental settings
In GML literature, there are two most popular settings for node property prediction regarding the relative sizes of train, validation, and test sets: one with a high label rate and one with a low label rate. In the high label rate setting, the train set encompasses 5050% of all graph nodes or more. This setting is common in heterophilous benchmarks, e.g., it is used by the datasets from Pei et al. (2020) and Platonov et al. (2023b), as well as by most datasets from OGB (Hu et al., 2020). In the low label rate setting, much smaller train set sizes are used (typically, no more than 1010% of all graph nodes). This setting is commonly used with the classic cora, citeseer, pubmed citation networks, and it is also used by the ogbn-products dataset from OGB. Both of these settings can appear in real-world GML usage scenarios, depending on the resources available for data labeling. It is important to provide predetermined splits to ensure experiments in different works are run in the same setting and their results are comparable, but it is also important to accommodate different needs of different research projects. Thus, for datasets in our benchmark we provide fixed splits for both settings. We refer to these splits as the RL (random low) and RH (random high) splits. Specifically, the RL split randomly divides nodes into train/validation/test sets with 1010%/1010%/8080% proportions, while the RH split randomly divides nodes into train/validation/test sets with 5050%/2525%/2525% proportions.
@@
The RL and RH data splits are random, as is common in current GML benchmarks. However, in real-world applications, data splits are often temporal, i.e., the labeled objects are the ones that appeared in the network earlier, while those that appeared later are not labeled and belong to the test set. Despite their prevalence in applications, temporal splits are very rarely used in current GML node property prediction benchmarks. To the best of our knowledge, the only such datasets with temporal splits available are citation networks from OGB, which represent only a single application. At the same time, temporal data splits may significantly affect the prediction problem and the model performance, as they often result in distributional shifts between train, validation, and test data. While some types of distributional shifts have been previously explored in the GML literature, e.g., shifts in node features (Gui et al., 2022) and shifts in graph structure (Bazhenov et al., 2023), realistic temporal distributional shifts often combine shifts in several aspects of data simultaneously (e.g., shifts in the distributions of node features, labels, and graph structural characteristics), and the effect of such realistic shifts on GML model performance is currently under-explored. To close this gap, we provide a temporal split for most datasets in our benchmark. We refer to this split as the TH (temporal high) split; it divides nodes into train/validation/test sets with 5050%/2525%/2525% proportions, i.e., exactly the same proportions as in the RH split, which allows comparing model results between the RH and TH splits to see how the complexity of the task changes when temporal distributional shifts are introduced.
@@
Further, many real-world networks are not static, but evolve over time. Thus, in many applications, not only are there no labels available for nodes that appear in the network later, but the nodes themselves (with their attributes and incident edges) are not available at training time. This setting is known in GML as the inductive setting. In contrast to the transductive setting in which the whole graph is available at training time (including the nodes for which predictions should be made), in the inductive setting validation and test nodes are not available at training time. Despite temporally evolving graphs being common in practical applications, most standard node property prediction datasets only provide the transductive setting. While it is well-known that GNNs, in contrast to some other GML methods like shallow node embeddings, can work not only in the transductive but also in the inductive setting (Hamilton et al., 2017), it is not well-explored how the lack of complete graph information at training time in the inductive setting affects GNN performance. Moreover, when the inductive setting is used for GNN evaluation in the current literature, the validation and test nodes are typically chosen randomly, which is not realistic, since in real-world applications the inductive setting is almost always induced by the temporal evolution of the graph. To fill this gap, for all datasets in our benchmark for which temporal information is available, we additionally provide the inductive experimental setting. We refer to this setting as THI (temporal high / inductive); it has the exact same data split as the transductive TH setting, but provides three snapshots of the graph: one for training, one for validation, and one for testing. This allows investigating how model performance changes between the transductive and the inductive setting. To the best of our knowledge, our work is the first to compare GNN performance between random and temporal data splits in the transductive setting under the same split ratios, and between transductive and inductive settings under the same temporal data split. This comparison allows us to investigate how much these differences can affect GNN performance.
@@ 5 Experiments
5.1 Models
Graph-agnostic baselines
The tasks that we consider in our benchmark, such as fraud detection and CTR prediction, are frequently tackled in industrial settings without considering the graph structure. Thus, as baselines, we use several models that do not have access to the graph structure. We call such models graph-agnostic. Comparing results between graph-agnostic and graph-aware models might reveal how much benefit using the graph structure brings for the considered task. First, as a simple baseline, we use ResMLP — an MLP augmented with skip-connections (He et al., 2016) and layer normalization (Ba et al., 2016). This model treats all nodes as independent data samples and does not use the graph structure. It has been shown that models of this type can serve as very strong baselines for industrial data with mixed numerical and categorical features (Gorishniy et al., 2021). Further, we consider GBDT models (Friedman, 2001). These models are very popular in industrial applications and are often considered to be better than neural networks at dealing with numerical features (Gorishniy et al., 2021, 2022) and regression tasks (Zhang et al., 2023), which makes them very relevant baselines for our benchmark. We use the three most popular GBDT implementations: XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018).
Further, we attempt to make originally graph-agnostic models stronger on our benchmark by providing them with some graph information through feature augmentation. Specifically, for each node in the graph, we aggregate features from all its one-hop neighbors in the graph, compute the mean, maximum, and minimum values of each feature, and append these statistics to the node’s original features. This mimics one step of GNN spatial graph convolution and allows graph-agnostic models to use some of the graph information that GNNs use. We call this feature augmentation strategy neighborhood feature aggregation (NFA) and denote models that use it with the -NFA suffix. A more detailed description of NFA is provided in Appendix B.
Graph neural networks
As our main models, we use four representative GNN architectures: two classic GNNs — GCN (Kipf & Welling, 2017) and GraphSAGE (Hamilton et al., 2017), and two attention-augmented GNNs — GAT (Veličković et al., 2018) and neighborhood-attention Graph Transformer (GT) (Shi et al., 2021). Note that GT is a local graph transformer with attention only over node neighborhoods, which is different from global graph transformers with all-to-all attention. For all GNNs, we use the modifications from Platonov et al. (2023b) that add skip connections, layer normalization, and MLP blocks to the models. We found that these modifications often significantly improve performance on our datasets.
It has been argued recently that in some GML datasets the graph structure does not actually help solve the considered tasks (Errica et al., 2020; Li et al., 2024b; Coupette et al., 2025; Bechler-Speicher et al., 2025). Thus, when new graph datasets are introduced, it is important to experimentally verify that their graph structure is actually helpful. For this, the performance of graph-agnostic and graph-aware models can be compared. To avoid other factors influencing the comparison, the two models being compared should be otherwise identical except that one has access to the graph structure and the other does not. Our set of models provides two such comparisons. First, graph-agnostic models can be compared with their NFA-augmented versions, which have access to graph neighborhood feature information. Second, ResMLP can be compared to GNNs: our GNNs have the exact same backbone that our ResMLP baseline has, except GNNs have spatial graph convolution modules added, which makes the comparison fair.
Graph foundation models
Recently, there has been a lot of interest in developing graph foundation models — models that can be applied to diverse graph datasets without or with minimal fine-tuning. However, we found that the current GFMs predominantly focus on text-attributed graphs and overlook the challenge of transferring a single model to datasets with different node feature sets. This prevents such models from being truly general, as graphs in real-world applications often have many different numerical and/or categorical node attributes that are important for solving the relevant tasks. We have found only a few GFMs that support node property prediction in graphs with arbitrary node attributes: SAMGPT (Yu et al., 2025), GCOPE (Zhao et al., 2024), OpenGraph (Xia et al., 2024), GraphFM (Lachi et al., 2024), AnyGraph (Xia & Huang, 2024), and TS-GNN (Finkelshtein et al., 2025). Of these models, only OpenGraph and AnyGraph have publicly available weights. Thus, we evaluate these two models. We commend the researchers who make their GFMs openly available and encourage others to do the same. For both models, we use the in-context learning (ICL) setting recommended by the authors for node classification, in which the task is cast as predicting links to virtual class nodes. Further, we were able to reproduce the pretraining of TS-GNN and GCOPE, and we also evaluate these two models (specifically, for TS-GNN we use the TS-Mean variant). We use the ICL setting for TS-GNN and the fine-tuning (FT) setting for GCOPE, as recommended by the authors (note that TS-GNN solves several least squares problems to fit its weights before making predictions, which can be considered a form of training, but, following the authors of this model, we still count it as ICL). However, neither of the considered models supports node regression; thus, we are only able to evaluate these models on our node classification datasets (while TS-GNN and GCOPE can support node regression in theory, their official implementations do not provide such support). Further, our experiments show that the considered GFMs cannot scale to large datasets.
5.2 Low label rate, random data split, transductive setting